OpenStudy (anonymous):

trying to solve system of differential equations using differential operator notation: x'' -2y'' -2y' = sin(t) x + y' = 0 I believe Diff. Operator notation is: D^2(x) - [(D^2 + 2)y] = sin(t) x + D(y) = 0 is this correct? if so, how do I set them up to solve by systematic elimination?

OpenStudy (anonymous):

D^2 - [(D^2 + 2D)y] = sin(t) x + Dy= 0

OpenStudy (anonymous):

osufan do you have the answer from the book?

OpenStudy (anonymous):

no, mark o, I don't have access to the book answer.